Circular No. 8432

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES AND RINGS OF SATURN C. C. Porco, CICLOPS, Space Science Institute, Boulder; and the

Cassini Imaging Science Team (cf. IAUC 8389) report the discovery of S/2004 S 5, a satellite orbiting in the region of Saturn III (and Tethys) and Saturn IV (Dione). The new satellite was seen in a series of six consecutive narrow-angle, clear-filter frames taken on Oct. 21, spanning 3 hr; it was also seen in two wide-angle images taken 12 days later. The spread of observations around the orbit is too short to report the orbital distance and configuration with confidence; if S/2004 S 5 is in a near-equatorial and near-circular orbit, it may be a coorbital satellite of Dione. S/2004 S 5 is ≈ 5 km across. Also, S/2004 S 3 (cf. IAUC 8401) was recovered at phase angle 64^{o} in a Cassini image taken on Oct. 17 (118 days after the discovery sequence of June 21). This puts the revolution period and semimajor axis at 0.62095 day and 140580 km. It is clear now that S/2004 S 3 is not the same object as S/2004 S 4 (IAUC 8401). Another object, S/2004 S 6, is also seen in the F-ring region, skirting the inner edge of the tenuous dust sheet that surrounds the F ring, in a sequence of movie images taken of that region at very high phase (153^{o}) over a period of 1 hr on Oct. 28; its approximate orbital period and semimajor axis are 0.61 day and 140000 km. The fact that the object was seen at 153^{o} phase may indicate that it is a clump with associated dust that is visible at high phase angles. It is possible, but not likely, that S/2004 S 6 is the same object as S/2004 S 4. More observations of this region, aimed at imaging orbiting satellites, are planned. A new narrow ring in the F-ring region, R/2004 S 2, is ≈ 300 km wide and is located at 138900 km between the ‘Atlas’ ring (R/2004 S 1; IAUC 8401) and the orbit of Saturn XVI (Prometheus).

SUPERNOVA 2004fc IN NGC 701 M. Salvo, B. Schmidt and S. Keller, Australian National University

(ANU), report that a spectrogram (range 400–700 nm) of SN 2004fc (cf. IAUC 8422), obtained with the ANU 2.3-m telescope (+ Double-Beam Spectrograph) at Siding Spring on Oct. 22.56 UT, shows it to be a type-II supernova before maximum light. The spectrum consists of a blue continuum with weak P-Cyg Balmer lines. Adopting the NASA/IPAC Extragalactic Database recession velocity of 1829 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 9700 km/s.

2004 November 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green